# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 4.2.1 Prior-focused regularization
In the medical domain, Baweja et al. (2018) are the first to explore EWC Kirkpatrick et al. (2017) to address the issue of catastrophic forgetting in neural networks when sequentially learning two distinct segmentation tasks (normal brain structure and white matter lesion). Specifically, the first task consists of multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the second task consists of segmentation of white matter lesions. The study demonstrates that EWC effectively reduces catastrophic forgetting in this challenging medical imaging context. Further, van Garderen et al. (2019) adopt EWC for glioma segmentation on different datasets in domain-shift arising from low and high-grade glioma in different datasets. In the context of histopathology breast cancer classification task, Chen and Tang (2022) showcase EWC capabilities for class incremental setting. McClure et al. (2018) introduce Distributed Weight Consolidation (DWC) as a continual learning method to consolidate weights of separate neural networks trained on independent datasets. Further, Zhang et al. (2023d) propose to compute the importance matrix jointly based on shape and semantics information in context for continual medical segmentation. The shape-based importance measures how sensitive a parameter is to shape properties in the images. and semantics-based importance measures a parameter’s sensitivity to reliable semantic predictions, ensuring that noisy or uncertain semantics do not influence the learning process. Then the selective regularization scheme is applied to penalize updates to model parameters with high joint importance weights. This ensures that critical shape and semantic knowledge from previous sites is not overwritten or forgotten.
Another prior-focused strategy is Synaptic Intelligence (SI) which alleviates catastrophic forgetting by allowing individual synapses (i.e., neurons) to estimate their importance for solving a learned task. Similarly to EWC, the approach penalizes changes to the most relevant synapses so that new tasks can be learned with minimal forgetting Zenke et al. (2017). A disadvantage is the need to distribute some extra parameters per weight in addition to their value, but in terms of data size, this is far less than providing the exemplars.
@@ 4.2.2 Data-focused regularization
Data-focused regularization method Li and Hoiem (2017) aims to distill knowledge from a model trained on the previous tasks to the model trained on the new task in order to consolidate previously learned knowledge. Typically, the previous model acts as a teacher and the current model as a student while adopting knowledge-distillation Gou et al. (2021) to avoid catastrophic forgetting. Usually, all the old data should be available in knowledge distillation which is not the case with CL; therefore, as an alternative, few old data or the current data or the generated data are explored for the same. The loss function has an additional distillation loss for replayed data. Each input is replayed with a soft target obtained using the stored model.
Silver and Mercer (2002) first proposed to use previous task model outputs given new task input images, mainly for improving new task performance. Later, in natural image classification application, Li and Hoiem (2017) re-introduced the concept as LwF Li and Hoiem (2017) method. A copy of the previous model parameters is stored before learning the new task and then it is used to get the soft labels for the new task as the target for the classifiers of the previous tasks. The available ground truth is used as the target for the new task classifier. LwF has been adopted in various medical applications like incremental brain MRI segmentation Ozdemir et al. (2018), chest x-ray classification Lenga et al. (2020), disease classification Derakhshani et al. (2022), histopathology tumor classification Kaustaban et al. (2022), etc.
Other works in natural image domain Jung et al. (2016); Zhang et al. (2020) have been introduced with LwF-related ideas, however, it has been shown that this strategy is vulnerable to domain shift between tasks Aljundi et al. (2017). It was pointed out that LwF would not result in good performance if the data distributions between different tasks are quite diverse Rannen et al. (2017). To overcome this, Rannen et al. (2017) facilitate incremental integration of shallow auto-encoders to constrain task features in their corresponding learned low dimensional space. Hence, they further trained an autoencoder for each task to learn the most important features corresponding to the task and used it to preserve knowledge.
@@
Karani et al. (2018)
MR brain segmentation: DIS (4 Ep.)
Domain specific batch normalization layer
Bayasi et al. (2021)
Skin lesion image classification: DIS (6 Ep.)
Architecture based (fixed but partitioned network)
Memmel et al. (2021)
Brain MRI Hippocampal Segmentation: DIS (2 Ep.)
Architecture (feature disentanglement through adversarial training)
González et al. (2022)
DIS with (2 Ep.) for MRI based Hippocampus Segmentation
Architecture based approach based on out of distribution detection concept: maintain multivariate Gaussians for all past learning created on batch normalization layer
@@ 4.3 Architecture-based
Architecture-based methods, also termed parameter isolation-based methods typically assign different parameters in a network to each task. This can be achieved by either fixing the architecture or dynamically extending the network Mallya and Lazebnik (2018); Hung et al. (2019); Mallya and Lazebnik (2018); Fernando et al. (2017); Yoon et al. (2017). Fixed architectures are limited by the network’s capacity, whereas dynamic architectures need more memory with every new task. Most of these methods use different parts of the network for each task which requires task identity during inference, but this identity information is usually unavailable.
Some works Yan et al. (2021a) aim to dynamically extend the feature extractor module by learning task-specific branches in it. Rusu et al. (2016) propose to use a dynamic architecture that blocks any changes to the network trained on previous knowledge and expands the architecture by allocating sub-networks with a fixed capacity to be trained with the new data and thus it keeps a pool of pre-trained models, one for each learned task. In contrast, Aljundi et al. (2017) propose a network of experts where each expert is a model trained given a specific task and a set of gating autoencoders that learn a representation for the task at hand, and, at inference time, automatically forward the test sample to the relevant expert. Another dynamic expandable model direction given by Yoon et al. (2017) expands the network using group sparse regularization to decide how many neurons to add at each layer and perform selective retraining. Mallya and Lazebnik (2018) propose to prune the weights with low magnitude and reuse them for the next task leaving the rest that are responsible for the previous tasks unchanged. In architecture-based approaches, typically the old knowledge is not exploited to learn the new, thus preventing knowledge transfer from a related task which is an important factor in continual learning.
@@
Bayasi et al. (2021) proposes a fixed-size architecture-based cl strategy where a subset of the network units is assigned to learn each domain separately. Their approach introduces a novel pruning criterion that allows a fixed network to learn new data domains sequentially over time. They identify culprit units associated with wrong classifications in each domain and use them for learning the new domain while freezing the non-culprit nodes.
@@
González et al. (2022) investigate the performance of continual learning methods in a task-agnostic setting, which better simulates dynamic clinical environments characterized by gradual population shifts. They propose an out-of-distribution detection-based solution that signals when to expand the model and select the best parameters during inference. Specifically, they learn a multivariate Gaussian on the last batch normalization layer of the deep model and store it in the memory. When there is out-of-distribution detection or domain shift is alerted based on a threshold on Mahalnobis distance from existing Gaussian in the memory and the new data, a new Gaussian is added to the memory. The closest Gaussian is used for inference and thus eliminates the need for domain ID during inference.
Karani et al. (2018) propose to use domain/task-specific batch norm layers to adapt to new MRI protocols while learning the segmentation of various brain regions. Entire batch-normalization layers were dedicated to modeling domain differences. However, this causes quadratic parameter increase with new tasks. As this approach dedicates specific batch-normalization parameters to each dataset/domain/task, task labels are necessary that determine to which dataset each sample belongs.
@@
Knowledge distillation + experience replay + multi-head
Ozdemir and Goksel (2019)
MRI based knee segmentation (2 Epi.) CIS
@@
Knowledge distillation + experience replay + multi-head
Ravishankar et al. (2019)
Chest x-ray view classification: DIS (2 Ep.), TIS (2 epi) Pneumothorax identification from X-ray: data-incremental (4 batches, each of 2K)
@@
Latent replay where samples are selected either randomly or importance based (farther from cluster centroids) + dynamic network (task specific feature transformer layers)
Lenga et al. (2020)
Chest x-ray classification: DIS (2 epi)
Comparative study: joint training, EWC, LwF
Li et al. (2020)
Dermoscopic images based skin disease classification: CIS (4-20 Ep.)
@@
Dual knowledge distillation (from old classifier (previous classes) and the fine-tuned classifier (a new FC layer fine-tuned on new classes)) + experience replay (fixed subset of data for each class is stored), during inference nearest-mean-of-exemplars method used for class prediction (Rebuffi et al. 2017) (where the mean is often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class)
Morgado et al. (2021)
DIS: dermatological imaging modality classification
@@
Comparative study: Regularization (EWC), rehearsal based (AGEM, experience replay)
Derakhshani et al. (2022)
TIS, CIS, DIS (4 epi)
Regularization, memory-replay (comparative: EWC, MAS, LwF, iCaRL, EEIL)
Ranem et al. (2022)
MRI based binary hippocampus segmentation: DIS (3 epi)
@@
Comparative study: transformers with EWC fisher matrix (Kirkpatrick et al. 2017), EWC with Riemannian walk (Chaudhry et al. 2018b), Modeling the Background (knowledge distillation and a modified Cross Entropy Loss) (Cermelli et al. 2020), Pseudo-labeling and Local Pod (multi-scale spatial distillation loss with pseudo labeling) (Douillard et al. 2021), Pooled Outputs Distillation (Douillard et al. 2020)
Liu et al. (2022)
@@
CIS: CT data based segmentation of abdomen organs (4 Ep. to sequentially include organs)
Rehearsal (latent replay) + regularization (knowledge distillation)
Akundi and Sivaswamy (2022)
Chest x-ray classification in CIS (5 epi)
@@
Regularization (knowledge distillation) + architecture (task specific head)
Shu et al. (2022)
DIS for fundus disease classification (2-4 epi)
@@
Replay (memory buffer optimization) + regularization (gradient regularization)
Kaustaban et al. (2022)
DIS, CIS, TIS, Data-IL
@@
Regularization, replay (comparative study: EWC, LwF, CoPE, iCaRL, A-GEM)
Chee et al. (2023)
CIS in classification setting for three applications: cancer classification, diabetic retinopathy classification, skin lesions classification
@@
Architecture (separate high-level descriptor for each task) + regularization (distillation loss) + replay (data based rehearsal, fixed memory per task)
Zhang et al. (2023b)
(a) CIS for multi-organ segmentation, (b) CIS for abdomen segmentation to liver tumor segmentation
@@
Architecture: organ specific segmentation head + pseudo labelling based knowledge distillation
Ji et al. (2023)
CIS for multi-organ segmentation
@@
Architecture: organ specific decoder + pseudo labelling based knowledge distillation
Zhang et al. (2023c)
CIS: continual disease classification in different image modality settings
@@
Experience replay + Architecture based (keeps task specific adapters and classifier heads)
Bai et al. (2023)
CIS
Regularization: knowledge distillation
Sadafi et al. (2023)
DIS, CIS, DIS+CIS
@@
Experience replay based rehearsal (representativeness of examples for each class and model uncertainty) + regularization (binary-cross entropy based distillation loss)
Yang et al. (2023b)
MedMNIST based disease classification: CIS+DIS (4 epi)
@@
Generative replay+ data-focused regularization
Bándi et al. (2023)
@@
DIS (cancer detection across organs such as breast, colon, and head-neck)
Comparative study of existing approaches: PackNet (architecture), EWC (regularization), GEM (rehearsal)
@@ 4.4 Hybrid techniques
Combining two or more of independent CL strategies has recently gained interest owing to the enhanced performance gained by harnessing the merits of different CL strategies. Ozdemir et al. (2018) mitigate catastrophic forgetting for incremental medical segmentation applications using a distillation loss inspired by LwF Li and Hoiem (2017). Further, they extend the approach by incorporating memory-replay strategies and task-specific segmentation heads. Specifically, representative images were selected based on abstraction layer response and content distance in the last embedding layer. Liu et al. (2022) also use knowledge distillation and a memory buffer to store the prototypical representation of different organ categories. Akundi and Sivaswamy (2022) also keep class-specific classification head along with knowledge distillation to avoid forgetting for chest x-ray classification application where data arrives in class incremental fashion. A pseudo-rehearsal technique was combined with task-specific dense layers for pneumothorax classification Ravishankar et al. (2019) by Ravishankar et al. (2019); however, it suffers from the linear increase of parameters and unbounded memory to store features for every domain.
@@
Li et al. (2020) introduce dual distillation as well as a fixed memory-based experience replay to continually learn the effective model. Specifically, they maintain an expert classifier which is nothing but the previous model fine-tuned on new classes after replacing the old fully-connected layer with a new one. Then a final classifier is distilled using the old classifier and the expert classifier. A fixed amount of old samples is also stored and replayed with current data while training the updated classifier. Then nearest-mean-of-exemplars (often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class) method is used for category prediction at the test time.
@@
Chee et al. (2023) utilize the dynamic expanding network, regularization, as well as data replay for continually learning new classes. A low-level feature extractor is shared across tasks but a high-level feature extractor is especially learned for each subsequent task. Further, an alternate training procedure so as to learn new classes, i.e., newly added high-level feature extractor (by freezing others) and learning the old classes (freezing the new task feature extractor) is followed. A fixed memory per task is kept to replay samples with current samples.
@@
Sadafi et al. (2023) use experience replay-based rehearsal and binary cross-entropy-based regularization terms for mitigating catastrophic forgetting. There is fixed memory and each class has an equal contribution to it. The sampling strategy for exemplar selection for a specific class is twofold: half of the samples are selected based on their distance from the mean; i.e, the closest samples to the class mean are selected, and another half of the samples for the class are selected based on model’s uncertainty for each sample as given in epistemic uncertainty estimation Mukhoti et al. (2021).
@@
Zhang et al. (2023b) exploit knowledge distillation as pseudo-labeling for old classes along with architecture-based CL strategy. In vanilla Swin UNETR architecture they replace the conventional output layer responsible for segmentation with organ-specific light-weight segmentation heads which thus enable independent predictions for any new or previously learned classes. Thus, there is a single encoder and decoder module and multiple organ-specific heads (few MLP layers) on the output of the encoder. In a similar fashion Ji et al. (2023) keep organ-specific encoder and apply knowledge distillation. They also apply network architecture search-based pruning on decoders to maintain network complexity. But in comparison to Zhang et al. (2023b) strategy the network complexity of maintaining multiple decoders is very high.
@@
Zhang et al. (2023c) propose a rehearsal and architecture-based approach. In order to effectively extract discriminative features from a pre-trained feature extractor for different diseases, a learnable lightweight adapter is added between consecutive convolutional stages at each subsequent episode. Furthermore, each task-specific classification head is also added at each step. Then in each episode model aim is to find optimal parameters in the newly added task-specific adapters and classification head using the training data from the new episode (new disease) and preserving a small subset for each previously learned disease.

### v3

@@ 4.2.1 Prior-focused regularization
In the medical domain, Baweja et al. (2018) are the first to explore EWC to address the issue of catastrophic forgetting in neural networks when sequentially learning two distinct segmentation tasks (normal brain structure and white matter lesion). Specifically, the first task consists of multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the second task consists of segmentation of white matter lesions. The study demonstrates that EWC effectively reduces catastrophic forgetting in this challenging medical imaging context. Further, van Garderen et al. (2019) adopt EWC for glioma segmentation on different datasets in domain-shift arising from low and high-grade glioma in different datasets. In the context of histopathology breast cancer classification task, Chen and Tang (2022) showcase EWC capabilities for class incremental setting. McClure et al. (2018) introduce Distributed Weight Consolidation (DWC) as a CL method to consolidate weights of separate neural networks trained on independent datasets. Further, Zhang et al. (2023d) propose to compute the importance matrix jointly based on shape and semantics information in context for continual medical segmentation. The shape-based importance measures how sensitive a parameter is to shape properties in the images and semantics-based importance measures a parameter’s sensitivity to reliable semantic predictions, ensuring that noisy or uncertain semantics do not influence the learning process. Then, the selective regularization scheme is applied to penalize updates to model parameters with high joint importance weights. This ensures that critical shape and semantic knowledge from previous sites is not overwritten or forgotten.
Another prior-focused strategy is Synaptic Intelligence (SI), which alleviates catastrophic forgetting by allowing individual synapses (i.e., neurons) to estimate their importance for solving a learned task. Similarly to EWC, the approach penalizes changes to the most relevant synapses so that new tasks can be learned with minimal forgetting Zenke et al. (2017). A disadvantage is the need to distribute some extra parameters per weight in addition to their value, but in terms of data size, this is far less than providing the exemplars.
@@ 4.2.2 Data-focused regularization
The data-focused regularization method Li and Hoiem (2017) aims to distill knowledge from a model trained on the previous tasks to the model trained on the new task in order to consolidate previously learned knowledge. Typically, the previous model acts as a teacher and the current model as a student while adopting knowledge-distillation Gou et al. (2021) to avoid catastrophic forgetting. Usually, all the old data should be available in knowledge distillation, which is not the case with CL; therefore, as an alternative, few old data or the current data or the generated data are explored for the same. The loss function has an additional distillation loss for replayed data. Each input is replayed with a soft target obtained using the stored model.
Silver and Mercer (2002) first proposed to use previous task model outputs given new task input images, mainly for improving new task performance. Later, in natural image classification application, Li and Hoiem (2017) re-introduced the concept as LwF technique. A copy of the previous model parameters is stored before learning the new task, and then it is used to get the soft labels for the new task as the target for the classifiers of the previous tasks. The available ground truth is used as the target for the new task classifier. LwF has been adopted in various medical applications like incremental brain MRI segmentation Ozdemir et al. (2018), chest x-ray classification Lenga et al. (2020), disease classification Derakhshani et al. (2022), histopathology tumor classification Kaustaban et al. (2022), etc.
Other works in the natural image processing domain Jung et al. (2016); Zhang et al. (2020) have been introduced with LwF-related ideas; however, it has been shown that this strategy is vulnerable to domain shift between tasks Aljundi et al. (2017). It was pointed out that LwF would not result in good performance if the data distributions between different tasks are quite diverse Rannen et al. (2017). To overcome this, Rannen et al. (2017) facilitate incremental integration of shallow auto-encoders to constrain task features in their corresponding learned low dimensional space. Hence, they further trained an autoencoder for each task to learn the most important features corresponding to the task and used it to preserve knowledge.
@@
Karani et al. (2018) MR brain segmentation: DIS (4 Ep.) Domain specific batch normalization layer
Bayasi et al. (2021) Skin lesion image classification: DIS (6 Ep.) Architecture based (fixed but partitioned network)
Memmel et al. (2021) Brain MRI Hippocampal Segmentation: DIS (2 Ep.) Architecture (feature disentanglement through adversarial training)
González et al. (2022) DIS with (2 Ep.) for MRI based Hippocampus Segmentation Architecture based approach based on out of distribution detection concept: maintain multivariate Gaussians for all past learning created on batch normalization layer
@@ 4.3 Architecture-based
Architecture-based methods, also termed parameter isolation-based methods typically assign different parameters in a network to each task. This can be achieved by either fixing the architecture or dynamically extending the network Mallya and Lazebnik (2018); Hung et al. (2019); Mallya and Lazebnik (2018); Fernando et al. (2017); Yoon et al. (2017). Fixed architectures are limited by the network’s capacity, whereas dynamic architectures need more memory with every new task. Most fixed architectures-based methods assign different parts of the network for each task, which requires task identity during inference, but this identity information is usually unavailable.
@@
In the fixed architecture category, Bayasi et al. (2021) proposes a cl strategy where a subset of the network units is assigned to learn each domain separately. Their approach introduces a novel pruning criterion that allows a fixed network to learn new data domains sequentially over time. They identify culprit units associated with wrong classifications in each domain and use them to learn the new domain while freezing the non-culprit nodes. Similarly, Mallya and Lazebnik (2018) propose to prune the weights with low magnitude and reuse them for the next task, while the remaining weights that are responsible for the previous tasks are kept unchanged. In architecture-based approaches, typically the old knowledge is not exploited to learn the new, thus preventing knowledge transfer from a related task which is an important factor in CL.
Some works aim to dynamically extend the feature extractor module by learning task-specific branches in it Yan et al. (2021a). Rusu et al. (2016) propose to use a dynamic architecture that blocks any changes to the network trained on previous knowledge and expands the architecture by allocating sub-networks with a fixed capacity to be trained with the new data, and thus it keeps a pool of pre-trained models, one for each learned task. In contrast, Aljundi et al. (2017) propose a network of experts where each expert is a model trained given a specific task and a set of gating autoencoders that learn a representation for the task at hand, and, at inference time, automatically forward the test sample to the relevant expert. Another dynamic expandable model direction given by Yoon et al. (2017) expands the network using group sparse regularization to decide how many neurons to add at each layer and perform selective retraining. Similarly, Karani et al. (2018) propose to use domain/task-specific batch norm layers to adapt to new MRI protocols while learning the segmentation of various brain regions. Entire batch-normalization layers were dedicated to modeling domain differences. However, this causes quadratic increase in parameters with new tasks. Moreover, as this approach dedicates specific batch-normalization parameters to each dataset/domain/task, task labels are necessary to determine to which dataset each sample belongs.
@@
In contrast, González et al. (2022) investigate the performance of CL methods in a task-agnostic setting, which better simulates dynamic clinical environments characterized by gradual population shifts. They propose an out-of-distribution detection-based solution that signals when to expand the model and select the best parameters during inference. Specifically, they learn a multivariate Gaussian on the last batch normalization layer of the deep model and store it in the memory. When there is out-of-distribution detection or domain shift is alerted based on a threshold on Mahalanobis distance from existing Gaussian in the memory and the new data, a new Gaussian is added to the memory. The closest Gaussian is used for inference and thus eliminates the need for domain ID during inference.
@@
Ozdemir et al. (2018) MRI-based Humerus and scapula segmentation: CIS (2 Ep.) Knowledge distillation + experience replay + multi-head
@@
Ozdemir and Goksel (2019) MRI based knee segmentation (2 Epi.) CIS Knowledge distillation + experience replay + multi-head
@@
Ravishankar et al. (2019) Chest x-ray view classification: DIS (2 Ep.), TIS (2 epi) Pneumothorax identification from X-ray: data-incremental (4 batches, each of 2K) Latent replay where samples are selected either randomly or importance based (farther from cluster centroids) + dynamic network (task specific feature transformer layers)
Lenga et al. (2020) Chest x-ray classification: DIS (2 epi) Comparative study: joint training, EWC, LwF
@@
Li et al. (2020) Dermoscopic images based skin disease classification: CIS (4-20 Ep.) Dual knowledge distillation (from old classifier (previous classes) and the fine-tuned classifier (a new FC layer fine-tuned on new classes)) + experience replay (fixed subset of data for each class is stored), during inference nearest-mean-of-exemplars method used for class prediction (Rebuffi et al. 2017) (where the mean is often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class)
@@
Morgado et al. (2021) DIS: dermatological imaging modality classification Comparative study: Regularization (EWC), rehearsal based (AGEM, experience replay)
Derakhshani et al. (2022) TIS, CIS, DIS (4 epi) Regularization, memory-replay (comparative: EWC, MAS, LwF, iCaRL, EEIL)
@@
Ranem et al. (2022) MRI based binary hippocampus segmentation: DIS (3 epi) Comparative study: transformers with EWC fisher matrix (Kirkpatrick et al. 2017), EWC with Riemannian walk (Chaudhry et al. 2018b), Modeling the Background (knowledge distillation and a modified Cross Entropy Loss) (Cermelli et al. 2020), Pseudo-labeling and Local Pod (multi-scale spatial distillation loss with pseudo labeling) (Douillard et al. 2021), Pooled Outputs Distillation (Douillard et al. 2020)
@@
Liu et al. (2022) CIS: CT data based segmentation of abdomen organs (4 Ep. to sequentially include organs) Rehearsal (latent replay) + regularization (knowledge distillation)
@@
Akundi and Sivaswamy (2022) Chest x-ray classification in CIS (5 epi) Regularization (knowledge distillation) + architecture (task specific head)
@@
Shu et al. (2022) DIS for fundus disease classification (2-4 epi) Replay (memory buffer optimization) + regularization (gradient regularization)
@@
Kaustaban et al. (2022) DIS, CIS, TIS, Data-IL Regularization, replay (comparative study: EWC, LwF, CoPE, iCaRL, A-GEM)
@@
Chee et al. (2023) CIS in classification setting for three applications: cancer classification, diabetic retinopathy classification, skin lesions classification Architecture (separate high-level descriptor for each task) + regularization (distillation loss) + replay (data based rehearsal, fixed memory per task)
@@
Zhang et al. (2023b) (a) CIS for multi-organ segmentation, (b) CIS for abdomen segmentation to liver tumor segmentation Architecture: organ specific segmentation head + pseudo labelling based knowledge distillation
@@
Ji et al. (2023) CIS for multi-organ segmentation Architecture: organ specific decoder + pseudo labelling based knowledge distillation
@@
Zhang et al. (2023c) CIS: continual disease classification in different image modality settings Experience replay + Architecture based (keeps task specific adapters and classifier heads)
Bai et al. (2023) CIS Regularization: knowledge distillation
@@
Sadafi et al. (2023) DIS, CIS, DIS+CIS Experience replay based rehearsal (representativeness of examples for each class and model uncertainty) + regularization (binary-cross entropy based distillation loss)
@@
Yang et al. (2023b) MedMNIST based disease classification: CIS+DIS (4 epi) Generative replay+ data-focused regularization
@@
Bándi et al. (2023) DIS (cancer detection across organs such as breast, colon, and head-neck) Comparative study of existing approaches: PackNet (architecture), EWC (regularization), GEM (rehearsal)
@@ 4.4 Hybrid techniques
Combining two or more of individual CL strategies has recently gained interest owing to the enhanced performance gained by harnessing the merits of different CL strategies. Ozdemir et al. (2018) mitigate catastrophic forgetting for incremental medical segmentation applications using a distillation loss inspired by LwF Li and Hoiem (2017). Further, they extend the approach by incorporating memory-replay strategies and task-specific segmentation heads. Specifically, representative images were selected based on abstraction layer response and content distance in the last embedding layer. Liu et al. (2022) also use knowledge distillation and a memory buffer to store the prototypical representation of different organ categories. Akundi and Sivaswamy (2022) also keep class-specific classification head along with knowledge distillation to avoid forgetting for chest x-ray classification application where data arrives in class incremental fashion. A pseudo-rehearsal technique was combined with task-specific dense layers for pneumothorax classification Ravishankar et al. (2019) by Ravishankar et al. (2019); however, it suffers from the linear increase of parameters and unbounded memory to store features for every domain.
@@
Li et al. (2020) introduce dual distillation as well as a fixed memory-based experience replay to continually learn the effective model. Specifically, they maintain an expert classifier, which is nothing but the previous model fine-tuned on new classes after replacing the old fully-connected layer with a new one. Then, a final classifier is distilled using the old classifier and the expert classifier. A fixed amount of old samples is also stored and replayed with current data while training the updated classifier. Then, the nearest-mean-of-exemplars (often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class) method is used for category prediction at the test time.
@@
Chee et al. (2023) utilize the dynamic expanding network, regularization, as well as data replay for continually learning new classes. A low-level feature extractor is shared across tasks, but a high-level feature extractor is especially learned for each subsequent task. Further, an alternate training procedure so as to learn new classes, i.e., newly added high-level feature extractor (by freezing others) and learning the old classes (freezing the new task feature extractor) is followed. A fixed memory per task is kept to replay samples with current samples.
@@
Sadafi et al. (2023) use experience replay-based rehearsal and binary cross-entropy-based regularization terms for mitigating catastrophic forgetting. There is fixed memory and each class has an equal contribution to it. The sampling strategy for exemplar selection for a specific class is twofold: half of the samples are selected based on their distance from the mean, i.e., the closest samples to the class mean are selected, and another half of the samples for the class are selected based on model’s uncertainty for each sample as given in epistemic uncertainty estimation Mukhoti et al. (2021).
@@
Zhang et al. (2023b) exploit knowledge distillation as pseudo-labeling for old classes along with architecture-based CL strategy. In vanilla Swin UNETR architecture, they replace the conventional output layer responsible for segmentation with organ-specific light-weight segmentation heads, which thus enable independent predictions for any new or previously learned classes. Thus, there is a single encoder and decoder module and multiple organ-specific heads (few MLP layers) on the output of the encoder. In a similar fashion Ji et al. (2023) keep organ-specific encoder and apply knowledge distillation. They also apply network architecture search-based pruning on decoders to maintain network complexity. But in comparison to Zhang et al. (2023b) strategy, the network complexity of maintaining multiple decoders is very high.
@@
Zhang et al. (2023c) propose a rehearsal and architecture-based approach. In order to effectively extract discriminative features from a pre-trained feature extractor for different diseases, a learnable lightweight adapter is added between consecutive convolutional stages at each subsequent task. Furthermore, each task-specific classification head is also added at each step. Then, in each task, the model aims to find optimal parameters in the newly added task-specific adapters and classification head using the training data from the new task (new disease) and preserve a small subset for each previously learned disease.